### v1

@@ 3 On $\Psi$-semismooth single-valued mappings
ad (iii): Since $\tilde{\Omega}\subset\Omega_{\Psi}\subset\Omega$ and the set $\Omega\setminus\tilde{\Omega}$ is negligible, so must be the set $\Omega\setminus\Omega_{\Psi}$ as well. For every $x\in\Omega_{\Psi}$ we deduce from (ii) that $\mathrm{conv}\,\overline{\nabla}F(\bar{x})$ is also a singleton and therefore $F$ is strictly differentiable at $x$ by Lemma 2.1.
Finally, the continuity of $\nabla F$ on $\Omega_{\Psi}$ is again a consequence of [28, Exercise 9.64]. ∎
The following theorem is a reformulation of [4, Theorem 3.2] using our notation. Consider the problem
$\min f(x)\quad\mbox{subject to}\quad x\in\omega,$
where $f:\mathbb{R}^{n}\to\mathbb{R}$ and $\omega\subset\mathbb{R}^{n}$ is a convex polyhedron.
@@
Assume that $f$ is semismooth with respect to some mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ and suppose that at each iteration of the BT-algorithm [33] instead of the oracle (1) one uses another
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $f(x)$ and one element $g(x)$ with $g(x)^{T}\in\Psi(x)$.
Then, if the sequence $z_{k}$ produced by the BT-algorithm remains bounded, there exists an accumulation point $\bar{z}$ of $z_{k}$ such that $0\in(\rm cocl\,\Psi)(\bar{z})+N_{\omega}(\bar{z})$.
Note that the condition $0\in(\rm cocl\,\Psi)(\bar{z})+N_{\omega}(\bar{z})$ is in fact a necessary optimality condition for a local minimizer because by Proposition 3.8 we have $\partial_{c}f(\bar{z})\subset(\rm cocl\,\Psi)(\bar{z})$.
4 On the SCD-ss property of graphically Lipschitzian mappings and consequences
@@
Consider $(x,y)\in\mathrm{gph}\,F\cap W$ and the corresponding pair $(u,f(u))=\Phi(x,y)$, where $W$ is as in Definition 2.16. Let $L\in{\mathcal{S}}F(x,y)$ and $A\in\overline{\nabla}f(u)$ such that $L=\nabla\Phi(x,y)^{-1}({\rm rge\;}(I,A)$ according to Lemma 2.17. Then
${\rm dist}((x-\bar{x},y-\bar{y}),L)$ $=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}\big{(}\nabla\Phi(x,y)(x-\bar{x},y-\bar{y})-(p,Ap)\big{)}\|$
$=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}\big{(}\Phi(x,y)-\Phi(\bar{x},\bar{y})+o(\|(x-\bar{x},y-\bar{y})\|)-(p,Ap)\big{)}\|$
$\leq\|\nabla\Phi(x,y)^{-1}\|\big{(}\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|+o(\|(x-\bar{x},y-\bar{y})\|)\big{)}$
$\leq\|\nabla\Phi(x,y)^{-1}\|\big{(}\|f(u)-f(\bar{u})-A(u-\bar{u})\|+o(\|(x-\bar{x},y-\bar{y})\|)\big{)}.$
@@
$\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|^{2}=\inf_{q\in\mathbb{R}^{n}}\|q\|^{2}+\|f(u)-f(\bar{u})-A(u-\bar{u})-Aq)\|^{2}$
$\geq\inf_{w\in\mathbb{R}^{n}}\frac{\|w\|^{2}}{\|A\|^{2}}+\|f(u)-f(\bar{u})-A(u-\bar{u})-w)\|^{2}=\frac{1}{1+\|A\|^{2}}\|f(u)-f(\bar{u})-A(u-\bar{u})\|^{2},$
@@
${\rm dist}((x-\bar{x},y-\bar{y}),L)=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}\big{(}\Phi(x,y)-\Phi(\bar{x},\bar{y})+o(\|(x-\bar{x},y-\bar{y})\|)-(p,Ap)\big{)}\|$
$\geq\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|-\|\nabla\Phi(x,y)^{-1}\|o(\|(x-\bar{x},y-\bar{y})\|)$
$\geq\frac{1}{\|\nabla\Phi(x,y)\|}\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|-\|\nabla\Phi(x,y)^{-1}\|o(\|(x-\bar{x},y-\bar{y})\|)$
$\geq\frac{1}{\|\nabla\Phi(x,y)\|\sqrt{1+\|A\|^{2}}}\|f(u)-f(\bar{u})-A(u-\bar{u})\|-\|\nabla\Phi(x,y)^{-1}\|o(\|(x-\bar{x},y-\bar{y})\|).$
@@
$\|u-\bar{u}\|\begin{cases}\geq\frac{\|(u-\bar{u},f(u)-f(\bar{u}))\|}{\sqrt{1+\kappa^{2}}}\geq\frac{\|(x-\bar{x},y-\bar{y})\|}{\|\nabla\Phi(x,y)^{-1}\|\sqrt{1+\kappa^{2}}}+o(\|(x-\bar{x},y-\bar{y})\|)&,\\ \leq\|(u-\bar{u},f(u)-f(\bar{u}))\|\leq\|\nabla\Phi(x,y)\|\|(x-\bar{x},y-\bar{y})\|+o(\|(x-\bar{x},y-\bar{y})\|),\end{cases}$
@@
$\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}\sup_{L\in{\mathcal{S}}F(x,y)}\frac{{\rm dist}((x-\bar{x},y-\bar{y}),L)}{\|(x-\bar{x},y-\bar{y})\|}=0\ \Longleftrightarrow\ \limsup_{u\to\bar{u}}\sup_{A\in\overline{\nabla}f(u)}\frac{\|f(u)-f(\bar{u})-A(u-\bar{u})\|}{\|u-\bar{u}\|}=0$
@@
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and is SCD-ss around $(\bar{x},\bar{y})$. Then there is an open neighborhood $W$ of $(\bar{x},\bar{y})$ such that for almost all $(x,y)\in\mathrm{gph}\,F\cap W$ (with respect to the $n$-dimensional Hausdorff measure) the mapping $F$ is strictly proto-differentiable at $(x,y)$.
@@
According to the notation of Definition 2.16, let $\Omega_{F}$ denote the set of all points in $\mathrm{gph}\,F\cap W$ where $F$ is strictly proto-differentiable and let $\Omega_{f}$ denote the set of all points $u\in U\subset\mathbb{R}^{n}$ where $f$ is strictly differentiable. We may also assume that $F$ is SCD-ss on $\mathrm{gph}\,F\cap W$ and therefore $f$ is ¯$\overline{\nabla}f$-ss on $U$ by Proposition 4.1. We conclude from Proposition 3.9 that the set $U\setminus\Omega_{f}$ has Lebesgue measure zero and the assertion follows from the relation
@@
which holds true by Theorem 2.18 and Lemmata 2.1, 2.17. ∎
@@ 5 On the semismooth property of solution maps to inclusions
$D^{*}S(x,y)(y^{*})\subset\Sigma(x,y)(y^{*}):=\{x^{*}\,\big{|}\,\exists z^{*}\in\mathbb{R}^{l}\mbox{ with }(x^{*},-y^{*})\in D^{*}F(x,y,0)(z^{*})\}.$ (9)
@@
We claim that there is some open neighborhood $U$ of $(\bar{x},\bar{y})$ and some real $\gamma>0$ such that for every pair $(x,y)\in\mathrm{gph}\,S\cap U$, every $z^{*}\in\mathbb{R}^{l}$, $\|z^{*}\|=1$ and every $(x^{*},y^{*})\in D^{*}F(x,y,0)(z^{*})$ one has $\|(x^{*},y^{*})\|\geq 1/\gamma$. Assume on the contrary that there are sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,S}}(\bar{x},\bar{y})$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in\mathrm{gph}\,D^{*}F(x_{k},y_{k},0)$ with $\|z_{k}^{*}\|=1$ and $\|(x_{k}^{*},y_{k}^{*})\|\to 0$ as $k$ tends to $\infty$. By possibly passing to a subsequence we can assume that $z_{k}^{*}$ converges to some $z^{*}$ with $\|z^{*}\|=1$ and by definition of the limiting coderivative we obtain $(0,0,z^{*})\in\mathrm{gph}\,D^{*}F(\bar{x},\bar{y},0)$ contradicting (8). Hence our claim holds true.
@@
and the inclusion (9) follows from [21, Theorem 4.46].
@@
$|\langle z^{*},0-0\rangle-\langle x^{*},x^{\prime}-x\rangle-\langle y^{*},y^{\prime}-y\rangle|\leq\epsilon\|(z^{*},x^{*},y^{*})\|\|(x^{\prime},y^{\prime},0)-(x,y,0)\|.$ (10)
Next consider $(y^{*},x^{*})\in\Gamma(x^{\prime},y^{\prime})=\mathrm{gph}\,\Sigma(x^{\prime},y^{\prime})$ and the corresponding $z^{*}$ verifying $(z^{*},x^{*},-y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},y^{\prime},0)$. By our claim we have $\|z^{*}\|\leq\gamma\|(x^{*},-y^{*})\|=\gamma\|(x^{*},y^{*})\|$ and from (10) we obtain
@@
This result extends slightly [21, Corollary 4.47], because estimate (9) holds true not only at the reference point $(\bar{x},\bar{y})$ but also on a neighborhood $U$ of it.
@@
$(x^{*},0)\in D^{*}F(\bar{x},\bar{y},0)(z^{*})\ \Rightarrow z^{*}=0,x^{*}=0$ (11)
holds. Then there is some open neighborhood $U$ of $\bar{x}$ such that $\sigma$ is Lipschitz continuous on $U$. If, in addition, $F$ is semismooth at $(x,\sigma(x),0)$ for some $x\in U$, then for every function $\varphi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$, which is semismooth at $(x,\sigma(x))$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$, the mapping $\theta:U\to\mathbb{R}$, $\theta(u):=\varphi(u,\sigma(u))$, is semismooth at $x$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
$\Psi(u):=\begin{cases}\{g_{x}^{T}+{x^{*}}^{T}\,\big{|}\,x^{*}\in\Sigma(u,\sigma(u))(g_{y}),(g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u))\}&\mbox{if $u\in U$,}\\ \emptyset&\mbox{else,}\end{cases}$
where $\Sigma$ is given by (9).

### v2

@@ 3 On $\Psi$-semismooth single-valued mappings
ad (iii): Since $\tilde{\Omega}\subset\Omega_{\Psi}\subset\Omega$ and the set $\Omega\setminus\tilde{\Omega}$ is negligible, so must be the set $\Omega\setminus\Omega_{\Psi}$ as well. For every $x\in\Omega_{\Psi}$ we deduce from (ii) that $\mathrm{conv}\,\overline{\nabla}F(x)$ is also a singleton and therefore $F$ is strictly differentiable at $x$ by Lemma 2.1. ∎
Corollary 3.9.
Let $F:\Omega\to\mathbb{R}^{m}$ be a mapping, where $\Omega\subset\mathbb{R}^{n}$ is open. Then the following two statements are equivalent and entail that $F$ is strictly differentiable on $\Omega$ up to a set of Lebesgue measure zero.
  1. 1.
    $F$ is semismooth on $\Omega$ with respect to some mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$.
  2. 2.
    $F$ is locally Lipschitz and $\mathrm{conv}\,\overline{\nabla}F$-ss on $\Omega$, i.e., G-semismooth.
Let us return to problem (1.1) under the additional assumption that $U_{\rm ad}$ is a convex polyhedron. The latter condition enables us to solve (1.1) numerically via the BT-algorithm which admits only $U_{\rm ad}$ given by affine linear inequalities and equalities.
Assume that $\theta$ is locally Lipschitz and suppose that at each iteration of the BT-algorithm instead of the oracle (1.2) one uses another
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $\theta(x)$ and one element $g(x)$ with $g(x)^{T}\in\Psi(x)$,
where $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ is a locally bounded mapping with $\mathrm{dom}\,\Psi$ being an open superset of $U_{\rm ad}$. Further assume that the condition
$\lim_{t\downarrow 0}\frac{\sup\{|\theta(x+td)-\theta(x)-t\langle g,d\rangle|\,\big{|}\,g\in\Psi(x+td)\}}{t}=0\ \forall x\in\mathbb{R}^{n}\ \forall d\in\mathbb{R}^{n}$ (3.11)
is fulfilled, which ensures that a certain line-search procedure in the BT-algorithm is finite. Then, as already observed by [8, Theorem 3.2], by replacing $\partial^{c}\theta$ with ${\rm cocl}\,\Psi$ in the convergence proof of the BT-algorithm, one can show the following result: If the sequence $z_{k}$ produced by the BT-algorithm remains bounded, we may conclude that there exists an accumulation point $\bar{z}$ of $z_{k}$ such that $0\in({\rm cocl}\,\Psi)(\bar{z})^{T}+N_{U_{\rm ad}}(\bar{z})$. Further, the above statement remains essentially true for other standard bundle methods.
If $\theta$ is $\Psi$-ss then condition (3.11) is clearly fulfilled and the condition $0\in({\rm cocl}\,\Psi)(\bar{z})+N_{U_{\rm ad}}(\bar{z})$ is in fact a necessary optimality condition for a local minimizer because, by Proposition 3.7, we have $\partial^{c}\theta(\bar{z})\subset({\rm cocl}\,\Psi)(\bar{z})$. We claim (without proof) that the latter assertion also holds true under the weaker condition (3.11). Further, condition (3.11) may be replaced by any other condition on $\Psi$ guaranteeing the finiteness of the line search procedure in order that the above convergence result of the BT-algorithm holds true. For instance, one could also use a condition related to upper semidifferentiability [1]. These weakened versions of semismoothness, however, will not be pursued in this paper because they do not satisfy a chain rule and are incompatible with the semismooth property of multifunctions.
Remark 3.10.
As already mentioned in the introduction, in the original versions of the BT-algorithm [36, 40] it is assumed that the objective $\theta$ is weakly semismooth, which may be characterized by directional differentiability of $\theta$ and the requirement that (3.11) holds with $\Psi=\partial^{c}\theta$.
4 On strict proto-differentiability and the SCD-ss property of graphically Lipschitzian mappings
At first we state the relationship between strict proto-differentiability of a mapping and the property that its SC derivative is a singleton.
@@ 4 On strict proto-differentiability and the SCD-ss property of graphically Lipschitzian mappings
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ has the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. If $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ and $\mathrm{gph}\,F$ is locally closed at $(\bar{x},\bar{y})$ then ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton. If, in addition, $F$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})$ then the reverse implication holds true as well.
Proof.
Assume that $F$ is strictly proto-differentiable, then we know by Lemma 2.4 that $\bar{L}:=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace. Further, from (2.6) we conclude that
Hence, $\bar{n}:=\dim\bar{L}\geq n$ and if $\bar{n}=n$ then ${\mathcal{S}}F(\bar{x},\bar{y})=\{\bar{L}\}$ follows. Now assume that $\bar{n}>n$ and choose an orthogonal basis $w_{i}$, $i=1,\ldots,\bar{n}$, of $\bar{L}$. Since $\bar{z}:=(\bar{x},\bar{y})\in(\mathrm{cl}\,{\cal O}_{F})\setminus{\cal O}_{F}$, there must be a sequence $z_{k}:=(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(\bar{x},\bar{y})$ satisfying $z_{k}\not=\bar{z}$ $\forall k$. Taking into account (2.6), for every $k$ sufficiently large and every $i=1,\ldots,\bar{n}$, we can find elements $w_{i}^{k}\in L_{k}:=T_{\mathrm{gph}\,F}(z_{k})$ with $\|w_{i}^{k}-w_{i}\|<1/\sqrt{\bar{n}}$. We claim now that for those $k$ the vectors $w_{i}^{k}$, $i=1,\ldots,\bar{n}$, are linearly independent. Assume on the contrary, that for some large $k$ these vectors are linearly dependent and we can find reals $\lambda_{i}^{k}$, $i=1,\ldots,\bar{n}$, with $\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}^{k}=0$ and $\sum_{i=1}^{\bar{n}}(\lambda_{i}^{k})^{2}=1.$ Then by orthogonality of $w_{i}$, $i=1,\ldots,\bar{n}$, together with the Cauchy-Schwarz inequality we obtain that
11 11 $=\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}\|\leq\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}w_{i}^{k}\|+\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(w_{i}-w_{i}^{k})\|$
$\leq\Big{(}\sum_{i=1}^{\bar{n}}(\lambda_{i}^{k})^{2}\Big{)}^{1/2}\Big{(}\sum_{i=1}^{\bar{n}}\|w_{i}-w_{i}^{k}\|^{2}\Big{)}^{1/2}<1\cdot(\sum_{i=1}^{\bar{n}}\frac{1}{\bar{n}})^{1/2}=1,$
a contradiction. Thus $w_{i}^{k}$, $i=1,\ldots,\bar{n}$, are linearly independent and $\dim L_{k}\geq\bar{n}$ follows. However, by definition of the set ${\cal O}_{F}$ we have $\dim L_{k}=n$, a contradiction. Hence $\bar{n}=n$ and ${\mathcal{S}}F(\bar{x},\bar{y})=\{\bar{L}\}$ holds true.
To show the second assertion, assume that $F$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})$ and let $W$, $\Phi$, $U$ and $f$ be as in Definition 2.10. Note that for all sequences $t_{k}\downarrow 0$, $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ and $(\hat{x}_{k},\hat{y}_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ such that $\lim_{k\to\infty}{((\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k}))}/{t_{k}}=:(\xi,\eta)$ exists, we have that $\Phi(x_{k},y_{k})=:(u_{k},f(u_{k})){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,f}}(\bar{u},f(\bar{u}))=\Phi(\bar{x},\bar{y})$, $\Phi(\hat{x}_{k},\hat{y}_{k})=:(\hat{u}_{k},f(\hat{u}_{k})){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,f}}(\bar{u},f(\bar{u}))$ and
$\lim_{k\to\infty}\frac{(\hat{u}_{k},f(\hat{u}_{k}))-(u_{k},f(u_{k}))}{t_{k}}$ $=\lim_{k\to\infty}\frac{\nabla\Phi(\bar{x},\bar{y})\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}+o(\|(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\|)}{t_{k}}$
$=\nabla\Phi(\bar{x},\bar{y})(\xi,\eta).$
For graphically Lipschitzian mappings, the SCD-ss property is closely related with the G-semismoothness.
@@
Consider $z:=(x,y)\in\mathrm{gph}\,F\cap W$ and the corresponding pair $(u,f(u))=\Phi(z)$, where $W$ is as in Definition 2.10. Let $L\in{\mathcal{S}}F(z)$ and $A\in\overline{\nabla}f(u)$ such that $L=\nabla\Phi(z)^{-1}({\rm rge\;}(I,A))$ according to Lemma 2.11. Then, setting $\bar{z}:=(\bar{x},\bar{y})$, we obtain that
${\rm dist}(z-\bar{z},L)$ $=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(z)^{-1}\big{(}\nabla\Phi(z)(z-\bar{z})-(p,Ap)\big{)}\|$
$=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(z)^{-1}\big{(}\Phi(z)-\Phi(\bar{z})+o(\|z-\bar{z}\|)-(p,Ap)\big{)}\|$
$\leq\|\nabla\Phi(z)^{-1}\|\big{(}\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|+o(\|z-\bar{z}\|)\big{)}$
$\leq\|\nabla\Phi(z)^{-1}\|\big{(}\|f(u)-f(\bar{u})-A(u-\bar{u})\|+o(\|z-\bar{z}\|)\big{)}.$
@@
$\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|^{2}=\inf_{q\in\mathbb{R}^{n}}\|q\|^{2}+\|f(u)-f(\bar{u})-A(u-\bar{u})-Aq\|^{2}$
$\geq\inf_{w\in\mathbb{R}^{n}}\frac{\|w\|^{2}}{\|A\|^{2}}+\|f(u)-f(\bar{u})-A(u-\bar{u})-w\|^{2}=\frac{1}{1+\|A\|^{2}}\|f(u)-f(\bar{u})-A(u-\bar{u})\|^{2},$
@@
${\rm dist}(z-\bar{z},L)$ $=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(z)^{-1}\big{(}\Phi(z)-\Phi(\bar{z})+o(\|z-\bar{z}\|)-(p,Ap)\big{)}\|$
$\geq\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(z)^{-1}(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|-\|\nabla\Phi(z)^{-1}\|o(\|z-\bar{z}\|)$
$\geq\frac{1}{\|\nabla\Phi(z)\|}\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|-\|\nabla\Phi(z)^{-1}\|o(\|z-\bar{z}\|)$
$\geq\frac{1}{\|\nabla\Phi(z)\|\sqrt{1+\|A\|^{2}}}\|f(u)-f(\bar{u})-A(u-\bar{u})\|-\|\nabla\Phi(z)^{-1}\|o(\|z-\bar{z}\|).$
@@
$\|u-\bar{u}\|\begin{cases}\geq\frac{\|(u-\bar{u},f(u)-f(\bar{u}))\|}{\sqrt{1+\kappa^{2}}}\geq\frac{\|z-\bar{z}\|}{\|\nabla\Phi(z)^{-1}\|\sqrt{1+\kappa^{2}}}+o(\|z-\bar{z}\|)&,\\ \leq\|(u-\bar{u},f(u)-f(\bar{u}))\|\leq\|\nabla\Phi(z)\|\|z-\bar{z}\|+o(\|z-\bar{z}\|),\end{cases}$
@@
$\mathop{{\rm Lim}\,{\rm sup}}_{z{\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}\bar{z}}\sup_{L\in{\mathcal{S}}F(z)}\frac{{\rm dist}(z-\bar{z},L)}{\|z-\bar{z}\|}=0\ \Leftrightarrow\ \limsup_{u\to\bar{u}}\sup_{A\in\overline{\nabla}f(u)}\frac{\|f(u)-f(\bar{u})-A(u-\bar{u})\|}{\|u-\bar{u}\|}=0$
@@
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and SCD-ss around $(\bar{x},\bar{y})$. Then there is an open neighborhood $W$ of $(\bar{x},\bar{y})$ such that for almost all $(x,y)\in\mathrm{gph}\,F\cap W$ (with respect to the $n$-dimensional Hausdorff measure) the mapping $F$ is strictly proto-differentiable at $(x,y)$ and ${\mathcal{S}}F(x,y)$ is a singleton.
@@
According to the notation of Definition 2.10, let $\Omega_{F}$ denote the set of all points in $\mathrm{gph}\,F\cap W$ where $F$ is strictly proto-differentiable and let $\Omega_{f}$ denote the set of all points $u\in U\subset\mathbb{R}^{n}$ where $f$ is strictly differentiable. We may also assume that $F$ is SCD-ss on $\mathrm{gph}\,F\cap W$ and therefore $f$ is ¯$\overline{\nabla}f$-ss on $U$ by Proposition 4.2. We conclude from Theorem 3.8 that the set $U\setminus\Omega_{f}$ has Lebesgue measure zero and the assertion follows from the relation
@@
which holds true by Theorem 4.1 and Lemmas 2.1, 2.11. ∎
@@ 5 On the semismooth property of solution maps to inclusions
$D^{*}S(x,y)(y^{*})\subset\Sigma(x,y)(y^{*}):=\{x^{*}\,\big{|}\,\exists z^{*}\in\mathbb{R}^{l}\mbox{ with }(x^{*},-y^{*})\in D^{*}F(x,y,0)(z^{*})\},\ y^{*}\in\mathbb{R}^{m}.$ (5.14)
@@
We claim that there is some open neighborhood $U$ of $(\bar{x},\bar{y})$ and some real $\gamma>0$ such that for every pair $(x,y)\in\mathrm{gph}\,S\cap U$, every $z^{*}\in\mathbb{R}^{l}$, $\|z^{*}\|=1$ and every $(x^{*},y^{*})\in D^{*}F(x,y,0)(z^{*})$ one has $\|(x^{*},y^{*})\|\geq 1/\gamma$. Assume on the contrary that there are sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,S}}(\bar{x},\bar{y})$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in\mathrm{gph}\,D^{*}F(x_{k},y_{k},0)$ with $\|z_{k}^{*}\|=1$ and $\|(x_{k}^{*},y_{k}^{*})\|\to 0$ as $k$ tends to $\infty$. By possibly passing to a subsequence we can assume that $z_{k}^{*}$ converges to some $z^{*}$ with $\|z^{*}\|=1$ and by definition of the limiting coderivative we obtain $(0,0,z^{*})\in\mathrm{gph}\,D^{*}F(\bar{x},\bar{y},0)$ contradicting (5.13). Hence our claim holds true.
@@
and the inclusion (5.14) follows from [28, Theorem 4.46].
@@
$|\langle z^{*},0-0\rangle-\langle x^{*},x^{\prime}-x\rangle-\langle y^{*},y^{\prime}-y\rangle|\leq\epsilon\|(x^{\prime},y^{\prime},0)-(x,y,0)\|\|(z^{*},x^{*},y^{*})\|.$ (5.15)
Next consider $(y^{*},x^{*})\in\Gamma(x^{\prime},y^{\prime})=\mathrm{gph}\,\Sigma(x^{\prime},y^{\prime})$ and the corresponding $z^{*}$ verifying $(z^{*},x^{*},-y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},y^{\prime},0)$. By our claim we have $\|z^{*}\|\leq\gamma\|(x^{*},-y^{*})\|=\gamma\|(x^{*},y^{*})\|$ and from (5.15) we obtain
@@
This result extends slightly [28, Corollary 4.47], because estimate (5.14) holds true not only at the reference point $(\bar{x},\bar{y})$ but also on a neighborhood $U$ of it.
@@
$(x^{*},0)\in D^{*}F(\bar{x},\bar{y},0)(z^{*})\ \Rightarrow z^{*}=0,\ x^{*}=0$ (5.16)
holds. Then there is some open neighborhood $U$ of $\bar{x}$ such that $\sigma$ is Lipschitz continuous on $U$. If, in addition, $F$ is semismooth at $(x,\sigma(x),0)$ for some $x\in U$, then for every function $\varphi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$, which is semismooth at $(x,\sigma(x))$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$, the mapping $\theta:U\to\mathbb{R}$, $\theta(u):=\varphi(u,\sigma(u))$, is semismooth at $x$ with respect to the mapping $\Theta:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
$\Theta(u):=\{g_{x}^{T}+{x^{*}}^{T}\,\big{|}\,x^{*}\in\Sigma(u,\sigma(u))(g_{y}),(g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u))\},\ u\in\mathrm{dom}\,\Theta:=U,$
where $\Sigma$ is given by (5.14).